OpenStudy (anonymous):

A movie theater charges 90 for adults and 70 for senior citizens. On a day when 325 people watched the movie, the total receipts were 24,950. How many among those who paid were adults? How many were seniors?

OpenStudy (anonymous):

a = adults s = seniors a + s = 325 90a + 70s = 24950 a = 325 - s 90(325 - s) + 70s = 24950 -20s + 29250 = 24950 20s = 4300 s = 215 a = 325 - 215 = 110

OpenStudy (anonymous):

Any questions?

OpenStudy (nincompoop):

systems of equation by substitution